# Spectral theory of the Neumann-Poincaré operator associated with multi-layer structures and analysis of plasmon mode splitting

### v2

@@ 5.2 An algebraic framework
In this subsection, we shall derive a precise connection between the plasmon mode and the choice of conductivities in the multi-layer confocal ellipses. To express the main results more clearly, we will defer the proofs to the subsequent section. By some straightforward but rather lengthy and tedious calculations, we can obtain the exact formulae to $|\mathbf{M}_{N,c}^{(n)}|$ or $|\mathbf{M}_{N,s}^{(n)}|$ for the structures with small number of layers, say $2\leqslant N\leqslant 5$. By observing these results, we induce the delicate result about the explicit formulae of $|\mathbf{M}_{N,c}^{(n)}|$ or $|\mathbf{M}_{N,s}^{(n)}|$ for all $N\in\mathbb{N}$ in the following theorem. For the convenience of description, we denote by $\mathbf{i}_{m}$ the multi-index $(i_{1},i_{2},\ldots,i_{m})$ and $\tau_{\mathbf{i}_{m}}$ its sign given by
@@
$f^{\pm}_{N}(\lambda)=\sum\limits_{k=0}^{\lfloor\frac{N}{2}\rfloor}\frac{\lambda^{N-2k}}{(2)^{2k}}\left(\sum\limits_{\mathbf{i}_{2k}\in C_{N}^{0,2k}}\tau_{\mathbf{i}_{2k}}\mathrm{e}^{2n\sum_{l=1}^{2k}(-1)^{l}\xi_{i_{l}}}\right)+O(\mathrm{e}^{-2n\tilde{\xi}}):=f_{N,\textup{r}}(\lambda)+O(\mathrm{e}^{-2n\tilde{\xi}}).$ (5.56)
@@
Remark 5.4.
We want to mention that even at plasmon mode, the plasmon resonance may still not be induced. We refer to [25] for details on inducing plasmon resonance. Here we show some explicit example based on Theorem 5.5. It is clearly that in the case of multi-layer thin strips in Theorem 5.5, the two possible plasmon modes are $\lambda=\pm\frac{1}{2}$. However, the plasmon resonance may not be induced for $\lambda=-\frac{1}{2}$. To better explain this, we consider a two-layered case and suppose that the background field $H=\cosh\xi\cos\eta$, i.e., $H$ is uniform background field. It then can be seen from (5.48) that
$\begin{split}u-H=&\cos\eta\frac{1}{e^{\xi}}\bm{e}^{T}\mathbb{F}^{(1)}_{2,l,c}\left(-\lambda\mathbf{I}_{2}-\left(\mathbf{K}_{2,c}^{(1)}\right)^{T}\right)^{-1}\tilde{\bm{s}}\\ =&\cos\eta\frac{\xi_{2}-\xi_{1}}{(\lambda-1/2)e^{\xi}}+o(\xi_{1}+\xi_{2})\quad\quad\mbox{in}\quad A_{0}.\end{split}$
It can be readily seen that the blow up on gradient of leading order term only occurs at $\lambda=1/2$ when $(\xi_{1}-\xi_{2})\sim(\xi_{1}+\xi_{2})$.
@@ 6.1 Eigenvalue problem
As mentioned above, the plasmon mode is parallel to the material configurations which make the parameter $-\lambda$ the eigenvalue of $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ or $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$ as $\tau\rightarrow 0$. So, we focus on the eigenvalues of the matrices $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, or the explicit formulae of $\left|\mathbf{M}_{N,c}^{(n)}\right|$ and $\left|\mathbf{M}_{N,c}^{(n)}\right|$ defined in (4.41) and (4.42), respectively, in this subsection. For this, we first define
@@
holds for all $N\geqslant 2$, $N\in\mathbb{N}$. ∎
@@
$\mathfrak{A}_{k}:=\sum_{\mathbf{i}_{k}\in C_{N-N_{0}+1}^{N_{0}-1,k}}\tau_{\mathbf{i}_{k}}\mathrm{e}^{2n\sum_{l=1}^{k}(-1)^{l}\xi_{i_{l}}},\quad\mathfrak{B}_{k}:=\sum_{\mathbf{i}_{k}\in C_{N-N_{0}}^{N_{0},k}}\tau_{\mathbf{i}_{k}}\mathrm{e}^{2n\sum_{l=1}^{k}(-1)^{l}\xi_{i_{l}}}.$ (6.67)
@@
Case ii $N$ is odd. The proof follows from a similar argument to the case that $N$ is even. We shall only briefly sketch it. First, it can be derived that
@@
In Proposition 3.1, we obtain that $\mathbb{K}_{N}^{*}$ can be symmetrizable and its spectrum lies in the interval $[-1/2,1/2]$. This, together with the fundamental theorem of algebra, implies that the polynomial $f^{\pm}_{N}(\lambda)$ has $N$ real roots that lies in the interval $[-1/2,1/2]$. By the definition (5.55) of $f_{N}^{\pm}(\lambda)$, it is easy to see that if $f_{N}^{+}(\lambda^{+})=0$, then $f_{N}^{-}(\lambda^{-})=0$. ∎
@@ 6.2 Asymptotic profiles of the plasmon modes
$=\sum\limits_{k=0}^{\lfloor\frac{N}{2}\rfloor}\frac{\lambda^{N-2k}}{(2)^{2k}}\left(\sum\limits_{\mathbf{i}_{2k}\in C_{N}^{0,2k}}\tau_{\mathbf{i}_{2k}}\mathrm{e}^{2n\sum_{l=1}^{2k}(-1)^{l}\xi_{i_{l}}}\right)+O(\mathrm{e}^{-2n\tilde{\xi}}).$
@@
Case ii $N$ is odd. From (6.77)–(6.79), it can be readily seen that
@@ 7 Numerical results
In this section, we present some numerical examples to corroborate our theoretical findings in the previous sections. From a theoretical point of view (see Theorems 5.25.3), in order to find the plasmon modes for any multi-layer structure, it suffices to find the roots of the characteristic polynomial $f^{\pm}_{N}(\lambda)$ defined in (5.55), namely, to find the eigenvalues of NP operator $\mathbb{K}_{N}^{*}$. Moreover, it is also important to understand the electric field distribution associated with each plasmon mode. We next investigate these scenarios numerically.
@@ 7.1 Mode splitting
Note that the ellipses are of the same focus. Define
$\lambda_{\pm}=(\lambda_{1}^{\pm},\lambda_{2}^{\pm},\ldots,\lambda_{N}^{\pm})$
where $\lambda_{k}^{\pm}$, $k=1,2,\ldots,N,$ is the root to $f_{N}^{\pm}(\lambda)$, respectively, we order them in descending way, that is, $\lambda_{1}^{\pm}>\lambda_{2}^{\pm}>\ldots>\lambda_{N}^{\pm}$. Table 1 shows the roots to the characteristic polynomials $f^{\pm}_{15}(\lambda)$ with $n=1$, respectively.
@@
$\sigma_{1}$4.5699 -4.2469 -3.7566 -3.1624 -2.5359 -1.9441 -1.4367 -1.0378
$\lambda_{+}$0.0725 -0.1468 -0.2078 -0.2540 -0.2868 -0.3081 - 0.3202
$\sigma_{1}$0.7467 -0.5461 -0.4128 -0.3262 -0.2710 -0.2374 -0.2193
@@
$\sigma_{1}$4.5605 -4.2123 -3.6894 -3.0657 -2.4225 -1.8312 -1.3391 -0.9636
$\lambda_{-}$0.0896 -0.1603 -0.2172 -0.2598 -0.2898 -0.3094 -0.3205
$\sigma_{1}$0.6960 -0.5144 -0.3943 -0.3162 -0.2662 -0.2355 -0.2188
@@
$\sigma_{1}$21.0821 -7.5514 -3.4990 -2.0042 -1.3877 -1.1258 -1.0274 -1.0027
$\lambda_{+}$0.0009 -0.0079 -0.0327 -0.0871 -0.1754 -0.2857 -0.3878 -0.4563
$\sigma_{1}$0.9966 -0.9690 -0.8771 -0.7033 -0.4807 -0.2728 -0.1263 -0.0457
@@
$\sigma_{1}$21.9003 -7.9149 -3.6661 -2.0803 -1.4219 -1.1401 -1.0320 -1.0034
$\lambda_{-}$0.0007 -0.0068 -0.0296 -0.0812 -0.1671 -0.2777 -0.3831 -0.4547
$\sigma_{1}$0.9973 -0.9733 -0.8882 -0.7206 -0.4989 -0.2858 -0.1324 -0.0474
@@
As previously discussed, concentric discs represent a special case where the radius of the confocal ellipse is sufficiently large (see Theorem 5.4). In order to verify this numerically, we let $\xi_{1}=L*N$ and $s=0.8$ in the setup (7.88). Set $N=17$ and $n=1$. Figure 7.4 shows that the norm of $|\lambda_{+}-\lambda_{-}|$ gradually decreases to zero with increasing $L$, which reveals the effect of rotational symmetry breaking on the splitting of the plasmon modes, and corroborates the theoretical results (5.56). We recover the result that the plasmon modes that split into even and odd parities in multi-layer confocal ellipses are degenerate into plasmon modes in multi-layer concentric discs in the limiting case $\tilde{\xi}\to\infty$.
@@ 7.2 Surface localization and geometrization of plasmon resonant modes
In subsection 7.1 above, we see that for the $N$-layer confocal ellipses and any fixed angular momentum $n\geqslant 1$, the $2N$ values of $\lambda\in[-1/2,1/2]$ such that there exists a non-trivial solution to (2.3). Then, the plasmon resonance occurs. Next, we seek to understand the distribution of the electric field associated with each plasmon mode for multi-layer structures based on Corollary 5.1. To this end, we first introduce the Drude model for modeling the parameter $\sigma_{1}$ with the operating frequency $\omega$
@@
We consider the multi-layer confocal ellipses with equidistant semi-major axis. For $N$-layer structure, set $b_{i}=R\cosh\xi_{i}$ satisfying
@@
Let $R=0.9$, $b_{N}=1$ ,and $T=0.2$. The eight plasmon resonant modes in the four-layer confocal ellipses designed by (7.90) are shown in Figure 7.5. We fix the scaling to be such that the electric field is normalized to the interval $[-1,1]$. The general splitting of the modes results from a breaking of rotational symmetry; however the modes specifically split into twice the number of layers because the multi-layer confocal ellipses still possesses two mirror planes. For the single-layer $m$-fold rotationally symmetric, we refer to [26].
For the multi-layer structures with rotationally symmetric geometry: concentric disks, the mode splitting in it is not owing to symmetry breaking. The reason for the mode splitting in it is owing to truncation of materials [24, 30], and becomes evident from the eight plasmon resonant modes in the eight-layer concentric disks in Figure 7.6. We can see that there are localized surface plasmons excited at almost each surface in multi-layer concentric disks. Because of the finite penetration depth of the light, some inner plasmon is not excited. There is thus a clear connection between the breaking of rotational symmetry in the $N$-layer confocal ellipses and plasmon hybridization in the $2N$-layer concentric disks.
@@ 8 Concluding remarks
In this paper, we developed a comprehensive mathematical framework to study electrostatics within multi-layer structures. We use layer potentials and symmetrization techniques to derive a perturbation formula in terms of the NP operator applicable to general multi-layer structures. We have characterized the spectral properties of the NP operator, revealing that the number of plasmon modes increases with the number of layers. The Fourier series-based matrix representation provides precise insights into plasmon mode behavior in complex geometries, such as multi-layer confocal ellipses. Our framework facilitates a deeper understanding of how material truncation and rotational symmetry breaking influence plasmon mode splitting, paving the way for the development of tailored metamaterial applications with enhanced functionalities. The computations of this paper do not apply to the dilation structure since different elliptic coordinates are required. We shall study this extension in our forthcoming works.

### v3

@@ 5.2 An algebraic framework
In this subsection, we shall derive a precise connection between the plasmon mode and the choice of conductivities in the multi-layered confocal ellipses. To express the main results more clearly, we will defer the proofs to the subsequent section. By some straightforward but rather lengthy and tedious calculations, we can obtain the exact formulae to $|\mathbf{M}_{N,c}^{(n)}|$ or $|\mathbf{M}_{N,s}^{(n)}|$ for the structures with small number of layers, say $2\leqslant N\leqslant 5$. By observing these results, we induce the delicate result about the explicit formulae of $|\mathbf{M}_{N,c}^{(n)}|$ or $|\mathbf{M}_{N,s}^{(n)}|$ for all $N\in\mathbb{N}$ in the following theorem. For the convenience of description, we denote by $\mathbf{i}_{m}$ the multi-index $(i_{1},i_{2},\ldots,i_{m})$ and $\tau_{\mathbf{i}_{m}}$ its sign given by
@@
[…]
@@ 6.1 Eigenvalue problem
As mentioned above, the plasmon mode is parallel to the material configurations which make the parameter $\lambda$ the eigenvalue of $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ or $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$ as $\tau\rightarrow 0$. So, we focus on the eigenvalues of the matrices $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, or the explicit formulae of $\left|\mathbf{M}_{N,c}^{(n)}\right|$ and $\left|\mathbf{M}_{N,c}^{(n)}\right|$ defined in (4.9) and (4.10), respectively, in this subsection. For this, we first define
@@
holds for all $\geqslant 2, NN\in\mathbb{N}$. ∎
@@
$\mathfrak{A}_{k}=\sum_{\mathbf{i}_{k}\in C_{N-N_{0}+1}^{N_{0}-1,k}}\tau_{\mathbf{i}_{k}}\mathrm{e}^{2n\sum_{l=1}^{k}(-1)^{l}\xi_{i_{l}}},\quad\mathfrak{B}_{k}=\sum_{\mathbf{i}_{k}\in C_{N-N_{0}}^{N_{0},k}}\tau_{\mathbf{i}_{k}}\mathrm{e}^{2n\sum_{l=1}^{k}(-1)^{l}\xi_{i_{l}}}.$ (6.9)
@@
Case $i$ NN is odd. The proof follows from a similar argument to the case that $N$ is even. We shall only briefly sketch it. First, it can be derived that
@@
In Proposition 3.1, we obtain that $\mathbb{K}_{N}^{*}$ can be symmetrizable and its spectrum lies in the interval [1/2,1/2][-1/2,1/2]. This, together with the fundamental theorem of algebra, implies that the polynomial $f^{\pm}_{N}(\lambda)$ has NN real roots that lies in the interval [1/2,1/2][-1/2,1/2]. By the definition (5.10) of $f_{N}^{\pm}(\lambda)$, it is easy to see that if $f_{N}^{+}(\lambda^{+})=0$, then $f_{N}^{-}(\lambda^{-})=0$. ∎
@@ 6.2 Asymptotic profiles of the plasmon modes
[…]
@@
Case $i$ NN is odd. From (6.19)–(6.21), it can be readily seen that
@@ 7 Numerical results
In this section, we present some numerical examples to corroborate our theoretical findings in the previous sections. From a theoretical point of view (see Theorems 5.25.3), in order to find the plasmon modes for any multi-layered structure, it suffices to find the roots of the characteristic polynomial $f^{\pm}_{N}(\lambda)$ defined in (5.10), namely, to find the eigenvalues of NP operator $\mathbb{K}_{N}^{*}$. Moreover, it is also important to understand the electric field distribution associated with each plasmon mode. We next investigate these scenarios numerically.
@@ 7.1 Mode splitting
Note that the ellipses are of the same focus. Define $\lambda_{\pm}=(\lambda_{1}^{\pm},\lambda_{2}^{\pm},\ldots,\lambda_{N}^{\pm}),$ where $\lambda_{k}^{\pm}$, $k=1,2,\ldots,N,$ is the root to $f_{N}^{\pm}(\lambda)$, respectively, we order them in descending way, that is, $\lambda_{1}^{\pm}>\lambda_{2}^{\pm}>\ldots>\lambda_{N}^{\pm}$. Table 1 shows the roots to the characteristic polynomials $f^{\pm}_{15}(\lambda)$ with $n=1$, respectively.
@@
$\sigma_{1}$ -4.5699 -4.2469 -3.7566 -3.1624 -2.5359 -1.9441 -1.4367 -1.0378
$\lambda_{+}$ -0.0725 -0.1468 -0.2078 -0.2540 -0.2868 -0.3081 - 0.3202
$\sigma_{1}$ -0.7467 -0.5461 -0.4128 -0.3262 -0.2710 -0.2374 -0.2193
@@
$\sigma_{1}$ -4.5605 -4.2123 -3.6894 -3.0657 -2.4225 -1.8312 -1.3391 -0.9636
$\lambda_{-}$ -0.0896 -0.1603 -0.2172 -0.2598 -0.2898 -0.3094 -0.3205
$\sigma_{1}$ -0.6960 -0.5144 -0.3943 -0.3162 -0.2662 -0.2355 -0.2188
@@
$\sigma_{1}$ -21.0821 -7.5514 -3.4990 -2.0042 -1.3877 -1.1258 -1.0274 -1.0027
$\lambda_{+}$ -0.0009 -0.0079 -0.0327 -0.0871 -0.1754 -0.2857 -0.3878 -0.4563
$\sigma_{1}$ -0.9966 -0.9690 -0.8771 -0.7033 -0.4807 -0.2728 -0.1263 -0.0457
@@
$\sigma_{1}$ -21.9003 -7.9149 -3.6661 -2.0803 -1.4219 -1.1401 -1.0320 -1.0034
$\lambda_{-}$ -0.0007 -0.0068 -0.0296 -0.0812 -0.1671 -0.2777 -0.3831 -0.4547
$\sigma_{1}$ -0.9973 -0.9733 -0.8882 -0.7206 -0.4989 -0.2858 -0.1324 -0.0474
@@
As previously discussed, concentric discs represent a special case where the radius of the confocal ellipse is sufficiently large (see Theorem 5.4). In order to verify this numerically, we let $\xi_{1}=L*N$ and $s=0.8$ in the setup (7.2). Set $N=17$ and $n=1$. Figure 7.4 shows that the norm of $|\lambda_{+}-\lambda_{-}|$ gradually decreases to zero with increasing $L$, which reveals the effect of rotational symmetry breaking on the splitting of the plasmon modes, and corroborates the theoretical results (5.11). We recover the result that the plasmon modes that split into even and odd parities in multi-layered confocal ellipses are degenerate into plasmon modes in multi-layered concentric discs in the limiting case $\tilde{\xi}\to\infty$.
@@ 7.2 Surface localization and geometrization of plasmon resonant modes
In subsection 7.1 above, we see that for the $N$-layer confocal ellipses and any fixed angular momentum $n\geqslant 1$, the $2N$ values of $\lambda\in[-1/2,1/2]$ such that there exists a non-trivial solution to (2.3). Then, the plasmon resonance occurs. Next, we seek to understand the distribution of the electric field associated with each plasmon mode for multi-layered structures based on Corollary 5.1. To this end, we first introduce the Drude model for modeling the parameter $\sigma_{1}$ with the operating frequency $\omega$
@@
We consider the multi-layered confocal ellipses with equidistant semi-major axis. For $N$-layer structure, set $b_{i}=R\cosh\xi_{i}$ satisfying
@@
Let $R=0.9$, $b_{N}=1$ ,and $T=0.2$. The eight plasmon resonant modes in the four-layer confocal ellipses designed by (7.4) are shown in Figure 7.5. We fix the scaling to be such that the electric field is normalized to the interval [1,1][-1,1]. The general splitting of the modes results from a breaking of rotational symmetry; however the modes specifically split into twice the number of layers because the multi-layered confocal ellipses still possesses two mirror planes. For the single-layer $m$-fold rotationally symmetric, we refer to [26].
For the multi-layered structures with rotationally symmetric geometry: concentric disks, the mode splitting in it is not owing to symmetry breaking. The reason for the mode splitting in it is owing to truncation of materials [24, 30], and becomes evident from the eight plasmon resonant modes in the eight-layer concentric disks in Figure 7.6. We can see that there are localized surface plasmons excited at almost each surface in multi-layered concentric disks. Because of the finite penetration depth of the light, some inner plasmon is not excited. There is thus a clear connection between the breaking of rotational symmetry in the $N$-layer confocal ellipses and plasmon hybridization in the $2N$-layer concentric disks.
@@ 8 Concluding remarks
In this paper, we developed a comprehensive mathematical framework to study electrostatics within multi-layered structures. We use layer potentials and symmetrization techniques to derive a perturbation formula in terms of the NP operator applicable to general multi-layered structures. We have characterized the spectral properties of the NP operator, revealing that the number of plasmon modes increases with the number of layers. The Fourier series-based matrix representation provides precise insights into plasmon mode behavior in complex geometries, such as multi-layered confocal ellipses. Our framework facilitates a deeper understanding of how material truncation and rotational symmetry breaking influence plasmon mode splitting, paving the way for the development of tailored metamaterial applications with enhanced functionalities. The computations of this paper do not apply to the dilation structure since different elliptic coordinates are required. We shall study this extension in our forthcoming works.